OpenStudy (anonymous):

how do i turn 72 over 420 into a mixed number

OpenStudy (unanimoose):

That can't be converted, it's not an improper fraction.

OpenStudy (mathstudent55):

For a fraction to be able to be converted into a mixed numeral, the fraction must have a numerator greater then the denominator. That means the fraction is greater than 1. In your case, the denominator is greater than the numerator. That means your fraction is less than 1, and it is a proper fraction, so it can't be turned into a mixed numeral.

OpenStudy (mathstudent55):

Now let's say your problem is 420 over 72. \(\dfrac{420}{72} \) Now we have an improper fraction, so we can turn it into a mixed numeral. You can start by reducing the fraction first. \(\dfrac{420}{72} = \dfrac{210}{36} = \dfrac{70}{12} = \dfrac{35}{6}\) Now we just divide 35 by 6 using long division, and we stop before going into decimal places. in other words, we want a remainder.  Now we see that 35 divided by 6 is 5 with remainder 5. That means \(\dfrac{420}{72} = 5~ \dfrac{5}{6}\)